[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Sep 29, 2017

Pablo Santos - NOAA Federal pablo.santos at noaa.gov
Fri Sep 29 12:37:38 UTC 2017


HSU:

Monica worked during the early morning hours on this problem and she found
the culprit behind the problems the evening and midnight shifts encountered
behind yesterday implemented RFC.

Things should work as normal for the 12Z cycle. Should you encounter any
problem, please contact us immediately.

Pablo

On Fri, Sep 29, 2017 at 5:20 AM, HSU Shift Log <nhc.tsbadmin at noaa.gov>
wrote:

> *HSU Shift Log for Fri Sep 29 2017 05:20:44 GMT-0400 (EDT)*
>
> *Shift:* A
>
> *Forecaster:* R Berg
>
> *Technical Issues Summary:*
>
> *ATCF issues:* Got red banners back after submitting guidance for Maria,
> Lee, and AL99. The trackers seemed to come back OK, although I did not
> receive SHIPS/LGEM. Also, the SDM reported that they did not receive the
> inputs for the HWRF/HMON submissions. Called Monica to look into the
> problems. It was eventually determined that the compute file timestamp
> needed to be updated and the file resubmitted to fix both problems.
> However, not sure why the timestamp didn't get updated appropriately on the
> first try.
>
> *N-AWIPS issues:*
>
> *AWIPS-2/NCP issues:*
>
> *Other Techincal issues:*
>
> *Miscellaneous issues:*




-- 
Dr. Pablo Santos
Acting Chief
Technology and Science Branch
National Hurricane Center, Miami, FL
305-229-4500
(c) 954-559-4965
http://www.hurricanes.gov
Facebook: *https://www.facebook.com/NWSNHC/
<https://www.facebook.com/NWSNHC/>*
Twitter: @NWSNHC @NHC_Atlantic @NHC_Pacific  @NHC_TAFB  @NHC_Surge
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170929/c2b6ee59/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list